# QED corrections to the correlated relativistic energy: one-photon processes

### v2

@@ Abstract
This work is a collection of initial calculations and formal considerations within the Salpeter–Sucher exact equal-time relativistic quantum electrodynamics framework. The calculations are carried out as preparation for the computation of pair, retardation, and radiative corrections to the relativistic energy of correlated two-spin-1/2-fermion systems. In this work, particular attention is paid to the retardation and the ‘one-loop’ self-energy corrections, which are known to be among the largest corrections to the correlated relativistic energy. The theoretical development is supplemented with identifying formal connections to the non-relativistic quantum electrodynamics framework, which is based on a correlated but non-relativistic reference, as well as to the ‘$1/Z$ approach’, which is built on a relativistic but one-particle zeroth order. The two complementary directions currently provide the theoretical framework for light atomic-molecular precision spectroscopy and heavy-atom phenomena. The present theoretical efforts pave the way for relativistic QED corrections to (explicitly) correlated relativistic computations.
@@ I Introduction
There are currently three distinct directions of research at the frontier of this field, (a) non-relativistic QED (nrQED): an $\alpha$ (and $\ln\alpha$) series expansion of the full theory about a correlated non-relativistic reference; (b) the ‘$1/Z$ approach’: an expansion of the full theory about an independent-particle relativistic reference; and (c) relativistic quantum chemistry: a practical combination of Dirac’s wave equation with the many-particle correlation techniques of quantum chemistry without first-principles account of relativistic QED in most applications.
@@ II Preliminaries to the Bethe–Salpeter equation: a formally exact wave equation for bound-state QED
including the $x=(\bm{r},t)$ space-time coordinate, the $\bm{p}_{a}=-\mathrm{i}\nabla_{a}$ momentum; $\bm{\alpha}_{a},\beta_{a}$ are the Dirac matrices, $1_{a}$ is the four-dimensional unit matrix, and $U_{a}$ labels the Coulomb potential of the fixed external charges (nuclei) interacting with the active particle of charge $z_{a}e$. Rationalized natural units, $\hbar=c=\varepsilon_{0}=1$, and $\alpha=e^{2}/(4\pi)$, are used throughout this work (except for Sec. IV and Appendices F and G, where this will be explicitly stated).
@@
Bound, stationary states (no time-dependent external fields being present) can be written in the form (Eq. (1.7) of Ref. 11 or Eqs. (2.20), (2.21) of Ref. 12)
@@
$S_{1}^{-1}(\bm{p}_{1},\varepsilon)S_{2}^{-1}(\bm{p}_{2},-\varepsilon)\psi(\bm{p}_{1},\bm{p}_{2},\varepsilon)=K\psi(\bm{p}_{1},\bm{p}_{2},\varepsilon)\ ,$ (11)
@@
The energy-momentum shift operators appearing in the interaction kernels, Eqs. (16) and (20), are extensively used throughout this work, so we collect here their effect on the one-particle propagators (see Eqs. (12), (13)):
@@
$S_{2}(\bm{p}_{2}-\bm{k},\omega-\varepsilon)=\eta_{2}(-\bm{k})\eta_{\epsilon}(\omega)S_{2}(\bm{p}_{2},-\varepsilon)\eta_{2}(\bm{k})\eta_{\epsilon}(-\omega)\ .$ (30)
@@ II.3 Perturbative account of $\cal{V}_{\text{QED}}$
Brillouin–Wigner perturbation theory is used in this context Sucher (1958); Douglas and Kroll (1974); Zhang (1996), since it is applicable even if the perturbation operator itself is energy-dependent,
@@
$E^{(1,0)}_{\varepsilon}$ $=\langle\Phi_{\text{np}}|\int_{-\infty}^{+\infty}\frac{\mathrm{d}\varepsilon}{-2\pi\mathrm{i}}(S^{(+)}_{1}+S^{(+)}_{2})K_{\Delta}(S_{1}+S_{2})(E-h_{1}-h_{2})^{-1}V_{\text{I}}L_{++}\Phi_{\text{np}}\rangle$
$\approx\langle\Phi_{\text{np}}|\int\frac{\mathrm{d}\varepsilon}{-2\pi\mathrm{i}}(S^{(+)}_{1}+S^{(+)}_{2})K_{\Delta}(S^{(+)}_{1}+S^{(+)}_{2})\Phi_{\text{np}}\rangle\;,$ (60)
@@ III.1 Single transverse-photon exchange
where $(1\leftrightarrow 2)$ denotes the same term with particle indices interchanged. If the zeroth-order solution, $(E_{\text{np}},\Phi_{\text{np}})$, was generated in the no-pair DC framework, then this result can be used without further manipulation. However, if the zeroth-order is the no-pair DCB solution, then $E_{\text{np}}$ already contains the instantaneous part of the transverse-photon exchange, i.e., the Breit term, so the perturbative correction must be computed only for the retarded part of the transverse correction to avoid double counting (Sec. III.2 and the discussion at the beginning of Sec. IV).
@@ III.2 Resummation for the instantaneous interaction ladder in the transverse-photon exchange correction
It is possible to formally add the full instantaneous (either Coulomb or Coulomb–Breit) ladder of positive- and negative-energy in-ladder states to the single transverse photon exchange (Fig. 2). In-ladder states mean the internal fermion lines between the rungs of the ladder. At the same time, the two in and out states of the ladder are forced to be of positive energy; this is trivially true when these states coincide with in/out states of the complete diagram (the external legs) because of the no-pair reference, but is a non-trivial restricition when they belong to internal lines connected to the emitting/absorbing vertex of the transverse photon. Besides derivational/computational convenience, the reason for this restriction is to treat all in/out states of the instantaneous ladder on the same footing.
@@ IV A unified treatment of transverse photon exchange and self-energy
In this section, atomic units are used, and factors of $c=1/\alpha$ are thus restored.
@@
$\bm{j}(\bm{k})=L_{++}\left[{z_{1}c}\eta_{1}(\bm{k})\tilde{\bm{\alpha}}_{1}(\bm{k})+{z_{2}c}\eta_{2}(\bm{k})\tilde{\bm{\alpha}}_{2}(\bm{k})\right]L_{++}\ ,$ (120)
@@ IV.1 Energy shift in the dipole approximation and a relativistic Bethe logarithm?
In the nrQED evaluation of the energy shift, Eq. (119) is split into high- and low-photon momentum parts. Different approximations are possible in the different momentum ranges; the parts of the energy shift are to be calculated using these approximations, with the obvious requirement that the final result cannot depend on the arbitrary separation of the energy ranges. The low-energy part of the shift is usually processed in the non-relativistic approximation (understood in the Foldy–Wouthuysen sense, with the exact non-relativistic energy eigenstate) combined with the dipole approximation $\eta_{a}(\bm{k})=\mathrm{e}^{\mathrm{i}\bm{k}\cdot\bm{r}_{a}}\approx 1$, and hence
$\bm{j}\approx\bm{j}_{0}=L_{++}\left[{z_{1}c}{\bm{\alpha}}_{1}+{z_{2}c}{\bm{\alpha}}_{2}\right]L_{++}\ .$ (123)
@@
$\langle\bm{j}_{0}\cdot\bm{j}_{0}\rangle=$ $z_{1}^{2}c^{2}\left\langle\bm{\alpha}_{1}L_{++}\bm{\alpha}_{1}\right\rangle+z_{2}^{2}c^{2}\left\langle\bm{\alpha}_{2}L_{++}\bm{\alpha}_{2}\right\rangle+2z_{1}z_{2}c^{2}\left\langle\bm{\alpha}_{1}L_{++}\bm{\alpha}_{2}\right\rangle$ (128)
@@ IV.3 Pilot numerical results and current limitations
${}^{\text{b}}$ Linear and logarithmic counterterms. The precise, regularized value of the non-relativistic counterterms is labelled with ${}^{\ast}$ and is $C_{\text{lin}}=6.125\ 587\ 7$, $C_{\text{log}}=45.501\ 047\ 1$ Korobov and Korobov (1999). Otherwise, we performed the direct evaluation of Eqs. (146)–(147) in the non-relativistic computation, and Eqs. (130)–(131) by Eqs. (144)–(145) in the relativistic case.
@@ V.1 On the asymptotic behaviour of the $k$-integrands, and perspectives on renormalization
The naive form of self-energy (Sec. III.3) contains a divergence associated with the electron mass being modified by its interaction with its own electromagnetic field. In order to obtain the finite, observable energy shift, the expressions must be renormalized by an appropriate subtraction of the divergent part. The degree of divergence depends on the approximations used in the evaluation of the self-energy. The investigation of the large $k$ limit of the self-energy correction is thus crucial to renormalization.
@@ Appendix A Some properties of the $\eta$ energy-momentum shift operator
and thus, the coordinate-space representation of $\eta$ is
@@ Appendix F A very brief overview of the variational DC/DCB method
$X=\begin{bmatrix}I&\emptyset&\emptyset&\emptyset\\ \emptyset&\frac{(\bm{\sigma}_{2}\cdot\bm{p}_{2})}{2m_{2}c}&\emptyset&\emptyset\\ \emptyset&\emptyset&\frac{(\bm{\sigma}_{1}\cdot\bm{p}_{1})}{2m_{1}c}&\emptyset\\ \emptyset&\emptyset&\emptyset&\frac{(\bm{\sigma}_{1}\cdot\bm{p}_{1})(\bm{\sigma}_{2}\cdot\bm{p}_{2})}{4m_{1}m_{2}c^{2}}\end{bmatrix}\ .$ (189)
We also note that factors of $c$ are temporarily restored; Hartree atomic units are most convenient for numerical computations. The basis functions are written as
@@ Appendix G Matrix elements of the velocity operator
For the computation of the relativistic Bethe logarithm and corresponding energy correction, Sec. IV.1, the matrix elements of the ‘kinetic balance transformed velocity operator’ $X\bm{j}X$ are required, see Eqs. (120) and (189). Following the conventions of previous work Jeszenszki et al. (2022),
$\bm{j}=cz_{1}\bm{\alpha}\boxtimes I+I\boxtimes cz_{2}\bm{\alpha}=\begin{bmatrix}\emptyset&cz_{2}\bm{\sigma}_{2}&cz_{1}\bm{\sigma}_{1}&\emptyset\\ cz_{2}\bm{\sigma}_{2}&\emptyset&\emptyset&cz_{1}\bm{\sigma}_{1}\\ cz_{1}\bm{\sigma}_{1}&\emptyset&\emptyset&cz_{2}\bm{\sigma}_{2}\\ \emptyset&cz_{1}\bm{\sigma}_{1}&cz_{2}\bm{\sigma}_{2}&\emptyset\end{bmatrix}\ ,$ (200)
@@
$X\bm{j}X=\begin{bmatrix}\emptyset&\frac{z_{2}\bm{\sigma}_{2}({\bm{\sigma}}_{2}\cdot\bm{p}_{2})}{2m_{2}}&\frac{z_{1}\bm{\sigma}_{1}({\bm{\sigma}}_{1}\cdot\bm{p}_{1})}{2m_{1}}&\emptyset\\ \frac{({\bm{\sigma}}_{2}\cdot\bm{p}_{2})z_{2}\bm{\sigma}_{2}}{2m_{2}}&\emptyset&\emptyset&\frac{p_{2}^{2}z_{1}\bm{\sigma}_{1}({\bm{\sigma}}_{1}\cdot\bm{p}_{1})}{8m_{1}m_{2}^{2}c^{2}}\\ \frac{({\bm{\sigma}}_{1}\cdot\bm{p}_{1})z_{1}\bm{\sigma}_{1}}{2m_{1}}&\emptyset&\emptyset&\frac{p_{1}^{2}z_{2}\bm{\sigma}_{2}({\bm{\sigma}}_{2}\cdot\bm{p}_{2})}{8m_{2}m_{1}^{2}c^{2}}\\ \emptyset&\frac{p_{2}^{2}({\bm{\sigma}}_{1}\cdot\bm{p}_{1})z_{1}\bm{\sigma}_{1}}{8m_{1}m_{2}^{2}c^{2}}&\frac{p_{1}^{2}({\bm{\sigma}}_{2}\cdot\bm{p}_{2})z_{2}\bm{\sigma}_{2}}{8m_{2}m_{1}^{2}c^{2}}&\emptyset\end{bmatrix}\ .$ (206)

### v3

@@ Abstract
This work is a collection of initial calculations and formal considerations within the Salpeter–Sucher exact equal-time relativistic quantum electrodynamics framework. The calculations are carried out as preparation for the computation of pair, retardation, and radiative corrections to the relativistic energy of correlated two-spin-1/2-fermion systems. In this work, particular attention is paid to the retardation and the ‘one-loop’ self-energy corrections, which are known to be among the largest corrections to the correlated relativistic energy. The theoretical development is supplemented with identifying formal connections to the non-relativistic quantum electrodynamics framework, which is based on a correlated but non-relativistic reference, as well as to the ‘$1/Z$ approach’, which is built on a relativistic but independent-particle zeroth order. The two complementary directions currently provide the theoretical framework for light atomic-molecular precision spectroscopy and heavy-atom phenomena. The present theoretical efforts pave the way for relativistic QED corrections to (explicitly) correlated relativistic computations.
@@ I Introduction
There are currently three distinct directions of research at the frontier of this field, (a) non-relativistic QED (nrQED) I. Eides et al. (2001); Jentschura and Adkins (2022); Yelkhovsky (2001); Korobov and Yelkhovsky (2001); Pachucki (2006); Yerokhin et al. (2021); Adkins et al. (2022): an $\alpha$ (and $\ln(\alpha)$) series expansion of the full theory about a correlated non-relativistic reference; (b) the ‘$1/Z$ approach’ Mohr et al. (1998); Shabaev (2002); Lindgren (2011); Yerokhin and Maiorova (2020); Artemyev et al. (2005); Kozhedub et al. (2019); Malyshev et al. (2019) an expansion of the full theory about an independent-particle relativistic reference; and (c) relativistic quantum chemistry Gorceix et al. (1987); Blundell et al. (1989); Parpia and Grant (1990); Dyall and Faegri, Jr. (2007); Reiher and Wolf (2015); Pyykkö (2012); Kutzelnigg (2012); Smits et al. (2023): a practical combination of Dirac’s wave equation with the many-particle correlation techniques of quantum chemistry without first-principles account of relativistic QED in most applications.
@@ II Preliminaries to the Bethe–Salpeter equation: a formally exact wave equation for bound-state QED
including the $x=(\bm{r},t)$ space-time coordinate, the $\bm{p}_{a}=-\mathrm{i}\nabla_{a}$ momentum; $\bm{\alpha}_{a},\beta_{a}$ are the Dirac matrices, $1_{a}$ is the four-dimensional unit matrix, and $U_{a}$ labels the Coulomb potential of the fixed external charges (nuclei) interacting with the active particle of charge $z_{a}e$. Rationalized natural units, $m=\hbar=c=\varepsilon_{0}=1$ and $\alpha=e^{2}/(4\pi)$ are used throughout this work.
@@
Bound, stationary states (no time-dependent external fields being present) can be written in the form (Eq. (1.7) of Ref. 33 or Eqs. (2.20), (2.21) of Ref. 34)
@@
$S_{1}^{-1}(\bm{p}_{1},\varepsilon)S_{2}^{-1}(\bm{p}_{2},-\varepsilon)\psi(\bm{p}_{1},\bm{p}_{2},\varepsilon)={\tilde{K}}\psi(\bm{p}_{1},\bm{p}_{2},\varepsilon)\ ,$ (11)
@@
The energy-momentum shift operators appearing in the interaction kernels, Eqs. (16) and (20), are extensively used throughout this work, so we collect here their effect on the one-particle propagators (Eqs. (12), (13)):
@@
$S_{2}(\bm{p}_{2}{+}\bm{k},\omega-\varepsilon)=\eta_{2}(-\bm{k})\eta_{\epsilon}(\omega)S_{2}(\bm{p}_{{}2},-\varepsilon)\eta_{2}(\bm{k})\eta_{\epsilon}(-\omega)\ .$ (30)
@@ II.3 Perturbative account of $\cal{V}_{\text{QED}}$
Brillouin–Wigner perturbation theory is used in this context Sucher (1958); Douglas and Kroll (1974); Zhang (1996) (see also Sec. 9.2 of Ref. 55 for an introduction), since it is applicable even if the perturbation operator itself is energy-dependent,
@@
$E^{(1,0)}_{\varepsilon}$ $=\langle\Phi_{\text{np}}|\int_{-\infty}^{+\infty}\frac{\mathrm{d}\varepsilon}{-2\pi\mathrm{i}}(S^{(+)}_{1}+S^{(+)}_{2})K_{\Delta}(S_{1}+S_{2})(E-h_{1}-h_{2})^{-1}V_{\text{I}}L_{++}{|}\Phi_{\text{np}}\rangle$
$\approx\langle\Phi_{\text{np}}|\int\frac{\mathrm{d}\varepsilon}{-2\pi\mathrm{i}}(S^{(+)}_{1}+S^{(+)}_{2})K_{\Delta}(S^{(+)}_{1}+S^{(+)}_{2}){|}\Phi_{\text{np}}\rangle\;,$ (60)
@@ III.1 Single transverse-photon exchange
where $(1\leftrightarrow 2)$ denotes the same term with particle indices interchanged (note that the $L_{++}$ factor could be neglected due to $\tilde{\bm{\alpha}}_{2}L_{++}\tilde{\bm{\alpha}}_{1}=L_{1+}\tilde{\bm{\alpha}}_{2}\tilde{\bm{\alpha}}_{1}L_{2+}$, but we keep it for later convenience). If the zeroth-order solution, $(E_{\text{np}},\Phi_{\text{np}})$, was generated in the no-pair DC framework, then this result can be used without further manipulation. However, if the zeroth-order is the no-pair DCB solution, then $E_{\text{np}}$ already contains the instantaneous part of the transverse-photon exchange, i.e., the Breit term, so the perturbative correction must be computed only for the retarded part of the transverse correction to avoid double counting (Sec. III.2 and the discussion at the beginning of Sec. IV).
@@ III.2 Resummation for the instantaneous interaction ladder in the transverse-photon exchange correction
It is possible to formally add the full instantaneous (either Coulomb or Coulomb–Breit) ladder of positive- and negative-energy in-ladder states to the single transverse photon exchange (Fig. 2). In-ladder states mean states represented by the internal fermion lines between the rungs of the ladder. At the same time, the two in and out states of the ladder are forced to be of positive energy; this is trivially true when these states coincide with in/out states of the complete diagram (the external legs) because of the no-pair reference, but is a non-trivial restricition when they belong to internal lines connected to the emitting/absorbing vertex of the transverse photon. Besides derivational/computational convenience, the reason for this restriction is to treat all in/out states of the instantaneous ladder on the same footing.
@@ IV A unified treatment of transverse photon exchange and self-energy
$\bm{j}(\bm{k})={\frac{1}{\alpha}}L_{++}\left[{z_{1}}\eta_{1}(\bm{k})\tilde{\bm{\alpha}}_{1}(\bm{k})+{z_{2}}\eta_{2}(\bm{k})\tilde{\bm{\alpha}}_{2}(\bm{k})\right]L_{++}\ ,$ (120)
@@
The factor of $1/\alpha$ was introduced in Eq. (120) to make the the leading ${\cal{O}}(m\alpha^{5})={\cal{O}}(\alpha^{3}E_{h})$ scaling of the retardation+self-energy correction manifest whenever an $\alpha$ expansion is meaningful (in the non-relativistic limit, particularly); see Appendix H for further details.
@@ IV.1 Energy shift in the dipole approximation and a relativistic Bethe logarithm?
In the nrQED evaluation of the energy shift, Eq. (119) is split into high- and low-photon momentum parts. Different approximations are possible in the different momentum ranges; the parts of the energy shift are to be calculated using these approximations, with the obvious requirement that the final result cannot depend on the arbitrary separation of the energy ranges. The low-energy part of the shift is usually processed in the non-relativistic approximation (understood in the Foldy–Wouthuysen sense, with the exact non-relativistic energy eigenstate) combined with the dipole approximation $\eta_{a}(\bm{k})=\mathrm{e}^{\mathrm{i}\bm{k}\cdot\bm{r}_{a}}\approx 1$, and hence $j_{i}(\bm{k})\approx\delta^{\perp}_{il}(\bm{k})(\bm{j}_{0})_{l}$, where
$\bm{j}_{0}={\frac{1}{\alpha}}L_{++}\left[{z_{1}}{\bm{\alpha}}_{1}+{z_{2}}{\bm{\alpha}}_{2}\right]L_{++}\ .$ (123)
@@
$\langle\bm{j}_{0}\cdot\bm{j}_{0}\rangle=$ ${\frac{1}{\alpha^{2}}\Big{[}}z_{1}^{2}\left\langle\bm{\alpha}_{1}L_{++}\bm{\alpha}_{1}\right\rangle+z_{2}^{2}\left\langle\bm{\alpha}_{2}L_{++}\bm{\alpha}_{2}\right\rangle+2z_{1}z_{2}\left\langle\bm{\alpha}_{1}L_{++}\bm{\alpha}_{2}\right\rangle{\Big{]}}$ (128)
@@ IV.3 Pilot numerical results and current limitations
${}^{\text{b}}$ Linear and logarithmic counterterms. The precise, regularized value of the non-relativistic counterterms is labelled with and is $C_{\text{lin}}=6.125\ 587\ 7$, $C_{\text{log}}=45.501\ 047\ 1$ Korobov and Korobov (1999). Otherwise, we performed the direct evaluation of Eqs. (146)–(147) in the non-relativistic computation, and Eqs. (130)–(131) by Eqs. (144)–(145) in the relativistic case.
@@ V.1 On the asymptotic behaviour of the $k$-integrands, and perspectives on renormalization
The naive form of self-energy (Sec. III.3) contains a divergence associated with the electron mass being modified by its interaction with its own electromagnetic field. In order to obtain the finite, observable energy shift, the expressions must be renormalized by an appropriate subtraction of the divergent part. The degree of divergence depends on the approximations used in the evaluation of the self-energy. The investigation of the large $k$ limit of the self-energy correction is thus crucial to renormalization. We will investigate the large $k$ asymptotics of the positive energy projected transverse and Coulomb self-energies (see Eqs. (118) and (104)), and -- to some extent -- the effect of negative energy parts (see Eqs. (LABEL:eq:SEnointTr) and (113)).
@@ Appendix A Some properties of the $\eta$ energy-momentum shift operator
and thus, the action of $\eta$ in coordinate-space representation is a simple multiplication:
@@ Appendix F A very brief overview of the variational DC/DCB method
$X=\begin{bmatrix}I&\emptyset&\emptyset&\emptyset\\ \emptyset&\frac{(\bm{\sigma}_{2}\cdot\bm{p}_{2})}{2m_{2}}&\emptyset&\emptyset\\ \emptyset&\emptyset&\frac{(\bm{\sigma}_{1}\cdot\bm{p}_{1})}{2m_{1}}&\emptyset\\ \emptyset&\emptyset&\emptyset&\frac{(\bm{\sigma}_{1}\cdot\bm{p}_{1})(\bm{\sigma}_{2}\cdot\bm{p}_{2})}{4m_{1}m_{2}}\end{bmatrix}\ .$ (187)
The basis functions are written as
@@ Appendix G Matrix elements of the velocity operator
For the computation of the relativistic Bethe logarithm and corresponding energy correction, Sec. IV.1, the matrix elements of the ‘kinetic balance transformed velocity operator’ $X\bm{j}X$ are required, Eqs. (120) and (187). Following the conventions of previous work Jeszenszki et al. (2022),
$\bm{j}={\frac{1}{\alpha}\Big{[}}z_{1}\bm{\alpha}\boxtimes I+I\boxtimes z_{2}\bm{\alpha}{\Big{]}}={\frac{1}{\alpha}}\begin{bmatrix}\emptyset&z_{2}\bm{\sigma}_{2}&z_{1}\bm{\sigma}_{1}&\emptyset\\ z_{2}\bm{\sigma}_{2}&\emptyset&\emptyset&z_{1}\bm{\sigma}_{1}\\ z_{1}\bm{\sigma}_{1}&\emptyset&\emptyset&z_{2}\bm{\sigma}_{2}\\ \emptyset&z_{1}\bm{\sigma}_{1}&z_{2}\bm{\sigma}_{2}&\emptyset\end{bmatrix}\ ,$ (194)
@@
$X\bm{j}X={\frac{1}{\alpha}}\begin{bmatrix}\emptyset&\frac{z_{2}\bm{\sigma}_{2}({\bm{\sigma}}_{2}\cdot\bm{p}_{2})}{2m_{2}}&\frac{z_{1}\bm{\sigma}_{1}({\bm{\sigma}}_{1}\cdot\bm{p}_{1})}{2m_{1}}&\emptyset\\ \frac{({\bm{\sigma}}_{2}\cdot\bm{p}_{2})z_{2}\bm{\sigma}_{2}}{2m_{2}}&\emptyset&\emptyset&\frac{p_{2}^{2}z_{1}\bm{\sigma}_{1}({\bm{\sigma}}_{1}\cdot\bm{p}_{1})}{8m_{1}m_{2}^{2}}\\ \frac{({\bm{\sigma}}_{1}\cdot\bm{p}_{1})z_{1}\bm{\sigma}_{1}}{2m_{1}}&\emptyset&\emptyset&\frac{p_{1}^{2}z_{2}\bm{\sigma}_{2}({\bm{\sigma}}_{2}\cdot\bm{p}_{2})}{8m_{2}m_{1}^{2}}\\ \emptyset&\frac{p_{2}^{2}({\bm{\sigma}}_{1}\cdot\bm{p}_{1})z_{1}\bm{\sigma}_{1}}{8m_{1}m_{2}^{2}}&\frac{p_{1}^{2}({\bm{\sigma}}_{2}\cdot\bm{p}_{2})z_{2}\bm{\sigma}_{2}}{8m_{2}m_{1}^{2}}&\emptyset\end{bmatrix}\ .$ (196)
@@
Appendix H On atomic units
Let us convert the last, retarded term of Eq. (119),
$\Delta{\cal{E}}=\frac{\alpha^{3}}{2\pi^{2}}\int\mathrm{d}^{3}\bm{k}\frac{1}{2{k}^{2}}\left\langle\bm{j}^{\dagger}(\bm{k})\left[\frac{E_{\text{DC}}-H_{\text{DC}}}{E_{\text{DC}}-H_{\text{DC}}-{{k}}}\right]\bm{j}(\bm{k})\right\rangle$ (202)
into atomic units $(m=e=\hbar=4\pi\varepsilon_{0}=1)$ by expressing everything in terms of the $a_{0}=1/(m\alpha)$ Bohr radius and the $E_{\text{h}}=m\alpha^{2}$ Hartree energy. This corresponds to the scaling
$\bm{r}=a_{0}\bm{r}^{\prime}\ \ ,\ \ \bm{k}=\frac{\alpha}{a_{0}}\bm{k}^{\prime}\ \ ,\ \ \bm{p}=\frac{1}{a_{0}}\bm{p}^{\prime}\ \ ,\ \ E_{\text{DC}}=E_{\text{h}}E^{\prime}_{\text{DC}}\ ,\ \ H_{\text{DC}}=E_{\text{h}}H^{\prime}_{\text{DC}}\ ,$ (203)
where primed quantities are dimensionless (the norm of the wave function is also rescaled). Note that we deliberately scale an extra $\alpha$ out of $\bm{k}$ (thereby, rescaling the integration variable), and thus, we formally have a factor of $E_{\text{h}}=\alpha/a_{0}$ in front of $\bm{k}^{\prime}$. Then, we find
$\Delta{\cal{E}}=E_{h}\frac{\alpha^{3}}{2\pi^{2}}\int\mathrm{d}^{3}\bm{k}^{\prime}\frac{1}{2{k^{\prime}}^{2}}\left\langle\bm{j}^{\dagger}(\alpha\bm{k}^{\prime})\left[\frac{E^{\prime}_{\text{DC}}-H^{\prime}_{\text{DC}}}{E^{\prime}_{\text{DC}}-H^{\prime}_{\text{DC}}-{{k}^{\prime}}}\right]\bm{j}(\alpha\bm{k}^{\prime})\right\rangle\ .$ (204)
The suppression of the exponent of $\eta_{a}$ by an extra factor of $\alpha$ once again motivates the dipole approximation $\eta_{a}\approx 1$, $j_{i}(\alpha\bm{k}^{\prime})\approx\delta^{\perp}_{il}(\bm{k}^{\prime})(\bm{j}_{0})_{l}$ (see Eq. (123)), leading to (after the same steps as in Sec. IV.1, and again discarding the logarithmically diverging part)
$\Delta E=-E_{\text{h}}\frac{2\alpha^{3}}{3\pi}\left\langle\bm{j}_{0}(H^{\prime}_{\text{DC}}-E^{\prime}_{\text{DC}})\bm{j}_{0}\right\rangle\ln(k_{0})\ .$ (205)
Of course, this transition to atomic units could have been performed directly after Eq. (133), but in this way, we can see the non-trivial $\alpha$-dependence of Eq. (204) when not resorting to the dipole approximation. By using Eqs. (203) and (196), we can see that the contribution of $\bm{j}_{0}$ to the expectation value is ${\cal{O}}(\alpha^{0})$, so the leading term in $\Delta E$ is indeed ${\cal{O}}(\alpha^{3}E_{h})={\cal{O}}(m\alpha^{5})$, as we know from nrQED.